OpenStudy (anonymous):

You roll a 6-sided die with faces numbered 1 through 6, and toss a coin. What is the probability of rolling a 2 or getting heads? A.1/6 B. 5/12 C. 2/3 D. 1/12

OpenStudy (anonymous):

\[P \left( A \cup B \right)=P \left( A \right)+P \left( B \right)\]

OpenStudy (anonymous):

Soo would it be 1/6 @surjithayer

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }+\frac{ 1 }{ 2 }=?\]